OpenStudy (anonymous):

Part 1: If you were solving a system of equations and you came to a statement like 1 = 3, what do you know about the solution(s) to this system? Part 2: Solve the following system and show all of your work. x − 2y = 14 x + 3y = 9

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

Part 1: No solution Part 2: Use cancellation.

OpenStudy (anonymous):

1. no solution 2. x - 2y = 14 --> x = 2y + 14 now sub 2y + 14 in for x in the 2nd equation to solve for y. Once y is found,sub that back into the 1st equation to solve for x.

OpenStudy (cwrw238):

part 1 there are no solutions part 2 use elimination: subtract equations to eliminate x: -2y - 3y = 14 - 9 -5y = 5 y = -1 now substitute y = -1 into the first equation x - 2(-1) = 14 and now its easy to find value of x

OpenStudy (anonymous):

The second problem can be solved with substitution or elimination, depending on which one you want to use

OpenStudy (anonymous):

well can do it so i can know how to do it

OpenStudy (anonymous):

thank you!!!!

OpenStudy (anonymous):

do you know how to do it ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

good deal :)